OpenStudy (anonymous):

A rectangle has a width of 2x-3cm and a length of 3x-4cm. Of the perimeter of the rectangle is 186cm, what is the value of x?

OpenStudy (anonymous):

this is optimization?

OpenStudy (anonymous):

is 20.4 one of your choices?

OpenStudy (anonymous):

No but there's a 20

OpenStudy (anonymous):

oh yes, i got it. it's 20cm

OpenStudy (anonymous):

186= 2(2x-3) +2(3x-4) 186=4x-6+6x-8 186=10x-14 200=10x x=20 make sense?

OpenStudy (anonymous):

if you want me to explain it to you, i can